OpenStudy (anonymous):

1/2+1/12 into a simplified fraction. steps would help to please.

OpenStudy (lalaly):

u need a common denominator 12 \[(1/2)*(6/6) +1/12 = 6/12 + 1/12 = 7/12\]

OpenStudy (anonymous):

Thank you so much for your help!

OpenStudy (lalaly):

:)